A child has a toy balloon with a volume of 2.80 liters. The


temperature of the balloon when it was filled was 25° C and the


pressure was 2.00 atm. If the child were to let go of the balloon


and it rose 3 kilometers into the sky where the pressure is


0.578 atm and the temperature is -9° C, what would the new


volume of the balloon be?

1
Expert's answer
2022-06-09T15:47:04-0400

Given:

P1 = 2.00 atm

V1 = 2.80 L

T1 = 25°C + 273.15 = 298.15 K

P2 = 0.578 atm

V2 = unknown

T2 = −9°C + 273.15 = 264.15 K


Formula: P1V1/T1 = P2V2/T2


Solution:

Combined gas law can be used.

Combined gas law can be expressed as:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (2.00 atm × 2.80 L × 264.15 K) / (0.578 atm × 298.15 K) = 8.58 L

V2 = 8.58 L


Answer: The new volume of the balloon would be 8.58 liters